OpenStudy (anonymous):

medal A nurse earns $37.50 an hour. For working the night shift, the nurse receives a 40% increase in pay. (a) What is the increase in hourly pay for working the night shift? $ ? (b) What is the hourly pay for working the night shift? $ ?

OpenStudy (anonymous):

she gets the original amount plus 40% 40% converted into decimal form is .40 So she gets the original amount + 40% of the original amount =37.50+37.50(.40) =37.50(1+.4) Factored out 37.50 =37.50(1.4) Added what was within the bracket =?